OpenStudy (anonymous):

If the legs of a right triangle are 16 and 12, would that make the hypotenuse 20?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

your opinion?

OpenStudy (anonymous):

I think it is true.

OpenStudy (loser66):

why?

OpenStudy (loser66):

Note: I didn't say it's right or wrong, just want you justify your thought

OpenStudy (anonymous):

I used a^2+b^2=c^2, so 16^2+12^2=c^2, 20 is what I came up with.

OpenStudy (loser66):

you got it. i am with you.

OpenStudy (loser66):

I am proud of you. my friend

OpenStudy (anonymous):

lol. Thanks again @Loser66 ! No way I would have known that without previous help from you. ( ͡° ͜ʖ ͡°)

OpenStudy (loser66):

glad to hear. hihihi